Bunson Burner has 3 liters of a 10% acid solution. How much pure acid must be added to produce a 50% acid solution?
Answer by stanbon(75887)

Bunson Burner has 3 liters of a 10% acid solution. How much pure acid must be added to produce a 50% acid solution?
-------------
Equation:
acid + acid = acid
-----
0.10*3 + 1*x = 0.50(3+x)
-------------
10*3 + 100x = 50*3 + 50x
50x = 40*3
-----
x = (4/5)*3 = 2.4 liters (amt. of pure acid needed)
---------------
Cheers,
Stan H.
-------------------